You invest $100 in a portfolio. The portfolio is composed of a risky asset with an expected rate of return of 12% and a standard deviation of 15% and a Treasury bill with a rate of return of 5%. What proportion of your total portfolio should be invested in the risky asset to form a portfolio with an expected rate of return of 9%?
Answer #1

Return for Risky Asset = Rr = 12%

Return on Treasury Bill = Rf = 5%

Let proportion invested in risky asset be wr = x

Proportion invested in Treasury bill = wf = 1 - x

=> Expected Return on Portfolio = wrRr + wfRf

Expected Return required = 9%

=> 9 = 12*x + 5*(1-x)

=> 9 = 12x + 5 - 5x

=> x = (9 - 5)/(12 - 5) = 0.5714 or 57.14%

Hence, 57.14% should be invested in Risky Asset